OpenStudy (milrawr):

e^4x-10=13 solve for x

OpenStudy (sunnnystrong):

\[e ^{4x}-10=13\] Any ideas @Milrawr ?

OpenStudy (milrawr):

the 10 is still in the exponent

OpenStudy (sunnnystrong):

oh okay \[e ^{4x-10}=13\]

OpenStudy (milrawr):

i'm not sure what to do i got 4x-10=e^13

OpenStudy (sunnnystrong):

Take the natural log of both sides ? :D

OpenStudy (milrawr):

i did that i just don't know what to do after that

OpenStudy (sunnnystrong):

Can i see some work?

OpenStudy (milrawr):

e^4x-10=13 lne(4x-10)=e^13 the lne cancel so 4x-10=e^13

OpenStudy (milrawr):

i'm probably doing it wrong :(

OpenStudy (sunnnystrong):

Hmm soo... Why would it be e^13? You would only do that to undo a natural log... So it's ln(13) Otherwise.. everything else looks right :)

OpenStudy (milrawr):

well i know the answer but it's not e^13

OpenStudy (milrawr):

sorry ln(13)

OpenStudy (sunnnystrong):

Yep! (:

OpenStudy (sunnnystrong):

so what is x?

OpenStudy (milrawr):

i don't know lol

OpenStudy (milrawr):

there's an answer key with it i just don't understand how to get the answer

OpenStudy (mathmale):

Milrawr: This discussion is taking a long while because of the ambiguity of your original expression. In a case like this, you could either take and share a screen shot of the original problem, or you could draw it in the Draw utility (below). Please draw the original function as carefully as you can.

OpenStudy (milrawr):

There's no drawing for this function. It's just an equation.

OpenStudy (mathmale):

No, Milrawr, I'm asking YOU to draw this function.

OpenStudy (mathmale):



OpenStudy (milrawr):

But there's no need to?? It's supposed to be done with algebra..

OpenStudy (mathmale):

Milrawr, your choice. You're missing the point here. This conversation has gone on and on primarily because the equation you shared is not precisely the same as the original function. Please decide: Either draw the original function accurately, or you're on your own.

OpenStudy (marcelie):

try this example .. its similar to your equation http://www.coolmath.com/sites/cmat/files/images/11-exponentials-06.gif

OpenStudy (milrawr):

i'm sorry that's dealing with growth rate it's not that bc there's an exact answer

OpenStudy (mathmale):

At this point, I'm leaving. Again and again I have asked you to share the original equation by drawing it, but you always counter that simple request. Good luck, Mil. Should you change your mind and try to cooperate, let me know.

OpenStudy (milrawr):

it's the first one sunny strong wrote and stop being so rude.

OpenStudy (mathmale):

As you wish, Milrawr. Good bye.

OpenStudy (milrawr):

lol i did it??

OpenStudy (mathmale):

Yes, you did it, Milrawr. You lost the chance to learn something valuable from someone who at first really wanted to help you.

OpenStudy (sshayer):

\[e ^{4x-10}=13\] taking natural log of both sides \[\ln e ^{4x-10}=\ln 13, (4x-10)\ln e=\ln 13\] 4x-10=ln 13=2.56 4x=2.56+10=12.56 x=?

OpenStudy (milrawr):

@mathmale I'm sorry I don't think the equation went through when I wrote it in the box. My apologies.